# Two-Stage Weighted Projection for Reliable Low-Complexity Cooperative and Non-Cooperative Localization

### v1

@@ V-A Positioning Performance
Fig. 3(a) and Fig. 3(b) present the positioning performance for the anchor-UE deployment depicted in Fig. 2, evaluated under both AWGN and multipath propagation scenarios. The results are obtained through extensive Monte Carlo simulations, where TOA errors are modeled in accordance with (40) and (44). As illustrated, in the AWGN scenario, our novel approach exhibits positioning accuracy superior to IPPM and NLS and matches the performance of WNLS. This observation is significant as WNLS is well-known for achieving CRLB in AWGN scenarios [3], a standard of optimality that is also met by our two-stage method. A similar trend is observed when extending the evaluation to multipath propagation conditions. Notably, even under multipath conditions, where NLOS effects and biases are prevalent, the performance of the proposed method remains remarkably close to the CRLB. This result also highlights that under favorable geometry and high SNR conditions, both our proposed approach and WNLS are asymptotically optimal.
@@
To further demonstrate the superior performance of TS-WPM, we investigated another challenging scenario characterized by low SNR conditions. For this evaluation, we reuse the anchor-UE deployment presented in Fig. 2, but instead of assuming a Tx power of 23 dBm, we reduce it to 13 dBm to induce low SNR conditions. All other simulation assumptions remain consistent with the previous assessment. Our evaluations indicated that, under AWGN conditions, TS-WPM performance closely matches that of WNLS; hence, we focus our discussion on the positioning performance observed under multipath propagation for low SNR scenarios. Fig. 6 illustrates the positioning performance for low SNR under multipath propagation. It is clearly evident that TS-WPM demonstrates superior performance compared to WNLS, in addition to IPPM and NLS. This superior performance is attributed to the amplified errors in TOA measurements at low SNRs, which are further exacerbated due to multipath propagation, where the NLOS bias offset depends on SNR and the underlying propagation conditions. To better understand this phenomenon, consider TOA estimation under multipath propagation for both high and low SNR scenarios. In high SNR conditions, the TOA estimation process is generally able to lock onto the first path or a path close to the first one, as these paths are distinct and detectable. Conversely, under low SNR conditions, the first path may become insignificant or undetectable due to poor SNR, leading to TOA estimates that are biased toward delayed paths. Although our evaluation framework does not explicitly estimate TOA, the TOA error modeling in (44) captures these effects by accounting for the prevailing SNR and propagation characteristics. Consequently, when TOA errors are significant–primarily due to a pronounced NLOS bias offset–the performance of WNLS deteriorates. In contrast, as described in Section II-A, the proposed approach incorporates a two-stage refinement process, which mitigates the impact of the NLOS bias offset by accurately estimating the range measurements associated with the reference anchor range and iteratively updating the UE position estimate and vice versa. As a result, the proposed approach ensures enhanced positioning even under low SNR conditions. Remarkably, even in high GDOP or low SNR scenarios, the proposed method achieves performance levels very close to the CRLB.
[…]
@@ V-B Computational Complexity
Having established the superior positioning performance of TS-WPM, we now analyze the computational complexity of the proposed TS-WPM and compare it with state-of-the-art algorithms, including NLS, WNLS, and IPPM. As outlined in Algorithm 1, each iteration of TS-WPM involves updating the UE location estimate, computing residuals, and refining the reference anchor range estimate across all $\small B$ anchors (see steps 1, 1, and 1). Each update step comprises basic vector norms and arithmetic operations, resulting in a per-iteration complexity of $\small O(B)$. Consequently, the overall complexity of TS-WPM is $\small O(N\cdot B)$, where $\small N$ represents the number of iterations required for convergence. Similarly, the computational complexity of IPPM is also $\small O(N\cdot B)$ [11, 10], which is identical to that of TS-WPM. In contrast, the computational complexity of NLS and WNLS is significantly higher due to additional matrix inversion and multiplication operations. Specifically, both methods exhibit a complexity of $\small O(N\cdot B\cdot D^{3})$ [3], where $\small D$ denotes the number of spatial dimensions of the UE coordinates. NLS and WNLS typically require fewer iterations than IPPM and TS-WPM, with TS-WPM necessitating additional iterations due to its two-stage refinement process. Our evaluations indicate that, on average, NLS and WNLS converge in 16-17 iterations, while IPPM and TS-WPM require more than 20 iterations. Despite its higher iteration count, TS-WPM remains computationally efficient compared to NLS and WNLS, as it avoids costly matrix inversions and minimizes the number of matrix multiplications.
@@ VI Concluding Remarks
-A Proof of Lemma 1
@@
This result is significant, as it shows that even under high GDOP conditions, the MSE of the proposed two-stage WPM remains directly proportional to the variance of pseudo-range measurements, which is not commonly observed in non-ideal settings [3]. By comparing (69) and (72), we observe that under high GDOP, $\small\lambda_{\min}\left(\mathbf{H}_{1}^{T}\mathbf{C}_{1}^{-1}\mathbf{H}_{1}\right)$ becomes insignificant, resulting in high MSE for the WNLS case. In contrast, the MSE with the novel TS-WPM in (72) trends similarly to low GDOP or favorable conditions, maintaining robust performance and reinforcing its effectiveness in non-ideal situations.

### v3

@@ V-A Positioning Performance
Fig. 3(a) and Fig. 3(b) present the positioning performance for the anchor-UE deployment depicted in Fig. 2, evaluated under both AWGN and multipath propagation scenarios. The results are obtained through extensive Monte Carlo simulations, where TOA errors are modeled in accordance with (41) and (45). As illustrated, in the AWGN scenario, our novel approach exhibits positioning accuracy superior to IPPM and NLS and matches the performance of WNLS. This observation is significant as WNLS is well-known for achieving CRLB in AWGN scenarios [3], a standard of optimality that is also met by our two-stage method. A similar trend is observed when extending the evaluation to multipath propagation conditions. Notably, even under multipath conditions, where NLOS effects and biases are prevalent, the performance of the proposed method remains remarkably close to the CRLB. This result also highlights that, under favorable geometry and high SNR conditions, both our proposed approach and WNLS are asymptotically optimal.
@@
To further demonstrate the superior performance of TS-WPM, we investigated another challenging scenario characterized by low SNR conditions. For this evaluation, we reuse the anchor-UE deployment presented in Fig. 2, but instead of assuming a Tx power of 23 dBm, we reduce it to 13 dBm to induce low SNR conditions. All other simulation assumptions remain consistent with the previous assessment. Our evaluations indicated that, under AWGN conditions, TS-WPM performance closely matches that of WNLS; hence, we focus our discussion on the positioning performance observed under multipath propagation for low SNR scenarios. Fig. 6 illustrates the positioning performance for low SNR under multipath propagation. It is clearly evident that TS-WPM demonstrates superior performance compared to WNLS, in addition to IPPM and NLS. This superior performance is attributed to the amplified errors in TOA measurements at low SNRs, which are further exacerbated due to multipath propagation, where the NLOS bias offset depends on SNR and the underlying propagation conditions. To better understand this phenomenon, consider TOA estimation under multipath propagation for both high and low SNR scenarios. In high SNR conditions, the TOA estimation process is generally able to lock onto the first path or a path close to the first one, as these paths are distinct and detectable. Conversely, under low SNR conditions, the first-arriving path may become insignificant or undetectable–not because only the first path is degraded, but because in multipath-rich environments (e.g., indoor settings), the first path is often weaker than subsequent reflected paths. Due to rich scattering, it may carry less energy and is more likely to fall below the detection threshold used in practical TOA estimation algorithms, such as correlation-based methods. In contrast, later-arriving stronger multipath components may still exceed the threshold and be detected, thereby biasing the TOA estimate toward delayed paths.
Although our evaluation framework does not explicitly estimate TOA, the TOA error modeling in (41) and (45) captures these effects by accounting for the prevailing SNR and propagation characteristics. Consequently, when TOA errors are significant–primarily due to a pronounced NLOS bias offset–the performance of WNLS deteriorates. In contrast, as described in Section II-A, the proposed approach incorporates a two-stage refinement process, which mitigates the impact of the NLOS bias offset by accurately estimating the range measurements associated with the reference anchor range and iteratively updating the UE position estimate and vice versa. As a result, the proposed approach ensures enhanced positioning even under low SNR conditions. Remarkably, even in high GDOP or low SNR scenarios, the proposed method achieves performance levels very close to the CRLB. Notably, the absolute magnitude of positioning errors is influenced by the system parameters used for evaluation, such as carrier frequency and bandwidth. While our evaluation adopts a representative 5G NR-compliant configuration (3.5 GHz carrier frequency with 5 MHz bandwidth) for demonstration purposes, the proposed TS-WPM is agnostic to these parameters and would yield significantly improved accuracy under wider bandwidths (e.g., 20-100 MHz) and lower carrier frequencies, as supported in 5G NR and beyond.
In our final evaluation, we assess the positioning performance of cooperative localization as outlined in Algorithm 2. For this assessment, we reuse the anchor-UE deployment depicted in Fig. 2, considering multipath propagation. However, unlike previous scenarios, we assume that each target UE has visibility to only two anchors, making collaborative measurements essential for accurate localization. The number of cooperative nodes, denoted by $\small N_{coop}$, is set to $\small[1,\,2,\,3]$, with nodes randomly selected from the UE drop. Fig. 7 illustrates the positioning performance of TS-WPM across different $\small N_{coop}$. The results clearly demonstrate the effectiveness of cooperative localization in enabling accurate positioning when an insufficient number of anchors are visible, with increasing the number of cooperative nodes leading to further accuracy improvements. Additionally, we compare the results with WNLS across different $\small N_{coop}$. It is crucial to note that WNLS, in its original form, fails to provide meaningful positioning estimates as the matrix in the normalization (see (72)) becomes ill-conditioned (non-invertible) due to highly unfavorable GDOP resulting from randomly selected collaborative nodes combined with limited anchor availability. This aligns with the insights established in Proposition 1, which show that WNLS results in high MSE in high GDOP scenarios. Addressing this, we employed a regularization factor to stabilize the matrix inversion process, a concept that is inherently incorporated in TS-WPM through a more systematic approach, as discussed in Section III-C. Our evaluation confirms that TS-WPM consistently outperforms WNLS due to its sophisticated normalization coupled with the two-stage approach, reinforcing its robustness under high GDOP conditions. Furthermore, we observed that cooperative localization with IPPM, under this setup, resulted in prohibitively high positioning errors, which is why we opted not to include a comparison with IPPM in this scenario. These findings underscore the effectiveness of TS-WPM in tackling challenging localization conditions where conventional approaches struggle to maintain accuracy.
@@ V-B Computational Complexity
Having established the superior positioning performance of TS-WPM, we now analyze the computational complexity of the proposed TS-WPM and compare it with state-of-the-art algorithms, including NLS, WNLS, and IPPM. As outlined in Algorithm 1, each iteration of TS-WPM involves updating the UE location estimate, computing residuals, and refining the reference anchor range estimate across all $\small B$ anchors (see steps 1, 1, and 1). Each update step comprises basic vector norms and arithmetic operations, resulting in a per-iteration complexity of $\small O(B\cdot D)$. Consequently, the overall complexity of TS-WPM is $\small O(N\cdot B\cdot D)$, where $\small N$ represents the number of iterations required for convergence and $\small D$ is the dimension of UE location coordinates (2 in our case). Similarly, the computational complexity of IPPM is also $\small O(N\cdot B\cdot D)$ [12, 11], which is identical to that of TS-WPM. In contrast, the computational complexity of NLS and WNLS is significantly higher due to additional matrix inversion and multiplication operations. Specifically, both methods exhibit a complexity of $\small O(N\cdot B\cdot D^{4})$ [3], where $\small D$ denotes the number of spatial dimensions of the UE coordinates. NLS and WNLS typically require fewer iterations than IPPM and TS-WPM, with TS-WPM necessitating additional iterations due to its two-stage refinement process. Our evaluations indicate that, on average, NLS and WNLS converge in 16-17 iterations, while IPPM and TS-WPM require more than 20 iterations. Despite its higher iteration count, TS-WPM remains computationally efficient compared to NLS and WNLS, as it avoids costly matrix inversions and minimizes the number of matrix multiplications.
Practical Considerations and Generalizability. While our evaluation framework assumes the availability of TOA/TDOA measurements, we acknowledge that in real-world deployments, TOA measurements are often subject to imperfections such as multipath-induced bias (NLOS bias), thermal noise, anchor synchronization errors, and residual UE clock offsets. To address these, our framework explicitly incorporates the effects of thermal noise and multipath through 3GPP-compliant CRLB formulations and the associated TOA error models, including a novel NLOS bias model that accounts for radio propagation characteristics and SNR, as described in Section IV. Moreover, the TDOA formulation inherently eliminates UE clock offset errors, assuming that the TOA measurements are acquired within a narrow measurement window (e.g., a few ms), during which the UE clock remains stable. While the current model assumes perfect synchronization among anchors, this is a valid assumption in many practical scenarios, such as 5G gNBs, Global Navigation Satellite System (GNSS)-synchronized UAV-mounted gNBs, or Low Earth Orbit (LEO) satellites equipped with GNSS receivers. Importantly, our framework is flexible enough to allow seamless integration of practical TOA estimation methods (e.g., correlation-based, Multiple Signal Classification (MUSIC), and Estimation of Signal Parameters via Rotational Invariant Techniques (ESPRIT))–as well as clock offset modeling and compensation mechanisms.
@@ VI Concluding Remarks
-A Key Definitions and Evaluation Metrics
Signal-to-Noise Ratio (SNR). Based on the received signal model in (29), the SNR between anchor $\small b$ and the target UE, denoted by $\small\gamma_{b}$, is defined as
$\small\displaystyle\gamma_{b}=\frac{P_{b}}{\sigma^{2}+NF_{l}},$ (46)
where $\small P_{b}$ represents the received signal power, $\small\sigma^{2}$ is variance of AWGN, and $\small NF_{l}$ denotes the noise figure in linear scale. To accurately compute $\small P_{b}$ and, consequently, $\small\gamma_{b}$, we adopt the dual-stripe channel model recommended in [24]. This model partitions the environment into line-of-sight (LOS) and NLOS regions, referred to as “stripes”. The stripe selection depends on the scenario–LOS stripes are used for outdoor UEs, while NLOS stripes are used for indoor UEs. Since our study assumes all UEs are located indoors, the NLOS stripe is employed to model large-scale fading.
The large-scale fading model accounts for distance-dependent path loss, shadowing, and penetration losses. The total path loss (in dB) is calculated as
$\small\displaystyle PL$ $\small\displaystyle=\max(15.3+37.6\log_{10}R,\,38.46+20\log_{10}R)$
$\small\displaystyle\quad 0.7d_{2D,\text{indoor}}+18.3\,n^{\left(\frac{n+2}{n+1}-0.46\right)}+L_{iw}+L_{ow},$ (47)
where $\small R$ is the 3D distance between the anchor and the UE, $\small d_{2D,\text{indoor}}$ is the horizontal indoor distance in meters, and $\small n$ is the number of floors penetrated, capturing the impact of buildings with multiple floors. $\small L_{iw}$ denotes the loss from internal walls (typically 5 dB per wall), and $\small L_{ow}$ accounts for penetration through external walls (typically 20 dB per wall).
Given the path loss, the SNR in dB is expressed as
$\small\displaystyle\gamma=P_{t}-PL-SF-N_{0}-NF,$ (48)
where $\small P_{t}$ is the transmit power in dBm (assumed to be 23 dBm, except for the low SNR case where it is reduced to 13 dBm), $\small PL$ is the path loss in dB as defined above, $\small SF$ is the shadow fading loss in dB, $\small NF$ is the noise figure in dB, and $\small N_{0}$ is the thermal noise power in dBm computed as
$\small\displaystyle N_{0}=-174+10\log_{10}(B),$ (49)
assuming a room temperature of $\small T=290$ K and a system bandwidth $\small B$ in Hz. The SNR computed using (48) is used throughout our evaluation to quantify signal quality for each anchor-UE link.
Geometric Dilution of Precision (GDOP). GDOP is a fundamental metric that quantifies the impact of anchor geometry on the accuracy of location estimation. Even when range measurements are accurate, unfavorable anchor distribution can amplify positioning errors due to high GDOP. Conversely, well-distributed anchors result in low GDOP and improved localization performance. To formalize this, consider a network of anchors where $\small\mathbf{r}$ denotes the true range measurements, $\small\boldsymbol{\theta}$ is the position parameter comprising target UE coordinates, and $\small\mathbf{H}$ is the corresponding Jacobian matrix relating the measurements to the UE location. The measurement model can be expressed as
$\small\displaystyle\mathbf{r}=\mathbf{H}\boldsymbol{\theta}.$ (50)
The GDOP is then defined by
$\small\displaystyle{\rm GDOP}=\sqrt{{\rm trace}\left(\left(\mathbf{H}^{T}\mathbf{H}\right)^{-1}\right)},$ (51)
where $\small(\cdot)^{T}$ denotes matrix transpose. In GPS settings [38], the matrix $\small(\mathbf{H}^{T}\mathbf{H})^{-1}$ takes the form
$\small\displaystyle\left(\mathbf{H}^{T}\mathbf{H}\right)^{-1}=\begin{bmatrix}D_{11}&D_{12}&D_{13}&D_{14}\\ D_{21}&D_{22}&D_{23}&D_{24}\\ D_{31}&D_{32}&D_{33}&D_{34}\\ D_{41}&D_{42}&D_{43}&D_{44}\end{bmatrix}$. (52)
from which various DOP metrics, including position DOP (PDOP), horizontal DOP (HDOP), vertical DOP (VDOP), and time DOP (TDOP) are derived.
$\small\displaystyle{\rm GDOP}$ $\small\displaystyle=\sqrt{D_{11}+D_{22}+D_{33}+D_{44}},$
$\small\displaystyle{\rm PDOP}$ $\small\displaystyle=\sqrt{D_{11}+D_{22}+D_{33}},$
$\small\displaystyle{\rm HDOP}$ $\small\displaystyle=\sqrt{D_{11}+D_{22}},{\rm VDOP}=\sqrt{D_{33}},\,{\rm TDOP}=\sqrt{D_{44}}.$
Here, PDOP quantifies the impact of anchor geometry on 3D positioning accuracy, HDOP on horizontal positioning accuracy, VDOP on vertical accuracy, and TDOP on the estimation of UE clock offset. GDOP reflects the overall geometric impact on all these components combined. For simplified scenarios with equal SNRs across anchors, the positioning error $\small P_{\text{err}}$ can be approximated as a function of PDOP and range error $\small r_{\text{err}}$ as
$\small\displaystyle P_{\text{err}}={\rm PDOP}\times r_{\text{err}}.$ (53)
In this article, we focus on 2D localization assuming no UE clock offset errors (or assuming TDOA inherently cancels out UE clock offset errors); hence, HDOP is of primary interest and is treated equivalently to GDOP and PDOP in our evaluation. To illustrate the impact of anchor geometry, we compute the GDOP for the anchor-UE deployments used in our experiments. Fig. 8 presents the CDF of GDOP under two distinct geometric settings: a favorable anchor deployment (see Fig. 2) and a non-ideal deployment (see Fig. 4). As evident from the results, good geometry results in significantly lower GDOP values compared to non-ideal geometry, which is consistent with the positioning results presented earlier–compare the low-GDOP results in Fig. 3 with the degraded accuracy under high-GDOP conditions shown in Fig. 5.
Figure 8: GDOP comparison: Good geometry vs. non-ideal geometry.
To contextualize the GDOP values observed in our experiments, Table II provides the qualitative interpretation of DOP values, as originally presented in [39]. A DOP value near 1 corresponds to ideal anchor geometry–such as when anchors are well spread–while values exceeding 10 typically reflect highly clustered or poorly distributed anchors, resulting in degraded localization accuracy. In our experiments, the GDOP values under favorable geometry consistently fall below 2, corresponding to the “Excellent” rating. For the non-ideal geometry, GDOP values span both “Excellent” and “Good” ranges. We intentionally avoided scenarios with GDOP in the “Fair” or “Poor” ranges, as such configurations would severely degrade the performance of benchmark methods such as WNLS, making the comparison unfair and less meaningful. Instead, by selecting challenging yet realistic configurations (i.e., “Good” vs. “Excellent” geometry), we demonstrate that the proposed TS-WPM achieves superior performance even under moderate and not so unfavorable conditions.
TABLE II: GDOP Rating Classification
GDOP Value Rating Description
1–2 Excellent
2–5 Good
5–10 Fair
[…]
-B Proof of Lemma 1
@@
This result is significant, as it shows that even under high GDOP conditions, the MSE of the proposed two-stage WPM remains directly proportional to the variance of pseudorange measurements, which is not commonly observed in non-ideal settings [3]. By comparing (78) and (81), we observe that under high GDOP, $\small\lambda_{\min}\left(\mathbf{H}_{1}^{T}\mathbf{C}_{1}^{-1}\mathbf{H}_{1}\right)$ becomes insignificant, resulting in high MSE for the WNLS case. In contrast, the MSE with the novel TS-WPM in (81) trends similarly to low GDOP or favorable conditions, maintaining robust performance and reinforcing its effectiveness in non-ideal situations.